Section 5

Differential Geometry and Global Analysis

In case of several authors, Invited Speakers are marked with a ^{∗}.

Dmitri Burago: Hard Balls Gas and Alexandrov Spaces of Curvature Bounded Above . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . II 289

Tobias H. Colding: Spaces with Ricci Curvature Bounds . . . . . . . . . . II 299
S. K. Donaldson: Lefschetz Fibrations in Symplectic Geometry …. II 309

Boris Dubrovin: Geometry and Analytic Theory of Frobenius Manifolds . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . II 315

Yakov Eliashberg: Invariants in Contact Topology . . . . . . . . . . . . . . . . II 327
S. Gallot: Curvature-Decreasing Maps are Volume-Decreasing . . . . . . II 339

Gerhard Huisken: Evolution of Hypersurfaces by Their Curvature in Riemannian Manifolds . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . II 349

Dominic Joyce: Compact Manifolds with Exceptional Holonomy …. II 361
François Labourie: Large Groups Actions on Manifolds . . . . . . . . . . . II 371
Joachim Lohkamp: Curvature Contents of Geometric Spaces . . . . . . . II 381

Franz Pedit and Ulrich Pinkall^{∗}: Quaternionic Analysis on Riemann Surfaces and Differential Geometry . . . . . . . . . . . . . . . . . . . . . . . II 389

Leonid Polterovich: Geometry on the Group of Hamiltonian Diffeomorphisms . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . II 401

Yongbin Ruan: Quantum Cohomology and its Application . . . . . . . . . II 411

Documenta Mathematica · Extra Volume ICM 1998 · II ·

288

Documenta Mathematica · Extra Volume ICM 1998 · II ·